OpenStudy (anonymous):

art A: Explain why the x-coordinates of the points where the graphs of the equations y = 8x and y = 2x + 2 intersect are the solutions of the equation 8x = 2x + 2. (4 points) Part B Make tables to find the solution to 8x = 2x + 2. Take the integer values of x between -3 and 3. (4 points) Part C: How can you solve the equation 8x = 2x + 2 graphically? (2 points)

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

y=8^x and y=2*x+2 all the x's are exponents in every part

OpenStudy (anonymous):

@itsbribro

OpenStudy (anonymous):

@Secret-Ninja @katkipe

OpenStudy (anonymous):

@thomaster @Preetha

OpenStudy (anonymous):

Dont mass-tag please

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

but i need help.